40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 2 19:05:13 2011, kiwamu, Update, IOO, peparation for low power MC : HWP+PBS installed attenuator.png
    Reply  Wed Aug 3 04:29:48 2011, kiwamu, Update, IOO, peparation for low power MC : HWP+PBS installed 
Message ID: 5103     Entry time: Wed Aug 3 04:29:48 2011     In reply to: 5097
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: peparation for low power MC : HWP+PBS installed 

[Suresh / Kiwamu]

 The incident beam power going into MC was decreased down to 20 mW by rotating the HWP that we set yesterday.

A 10% beam splitter which was sitting before MCREFL_PD was replaced by a perfect reflector so that all the power goes into the PD.

And we confirmed that MC can be still locked by increasing C1IOO-MC_REFL_GAIN. Some modifications in the Autolocker script need to be done later.

Also we opened the aperture of the MC2F camera to clearly see the low power beam spot.

 

WE ARE READY FOR THE VENT !!

Power after the EOM = 1.27 W

Power after the HWPs and PBS = 20.2 mW

Power on MCREFL = 20 mW (MC unlocked)

MCREFL_DC = 0.66 V (with MC locked)

 

Quote:

After we finish the measurement we will go to the low power state.

 

ELOG V3.1.3-